OpenStudy (anonymous):

AP Calculus: Find the intercepts: y=x^3-x^2-12x

OpenStudy (anonymous):

So far i found the derivative, y=x^3-x^2-12x

hartnn (hartnn):

To get x-intercept put y=0 in the equation to get y-intercept put x=0 in the equation

OpenStudy (anonymous):

goes beyond that. It has to be factored first

OpenStudy (anonymous):

@hartnn the question is asking to find the intercepts

OpenStudy (anonymous):

how do i factor it

hartnn (hartnn):

why do u need derivative to find intercepts ?

OpenStudy (anonymous):

oh so im doing it wrong/

OpenStudy (anonymous):

\[x^3-x^2-12x\]\[x(x^2-x-12)\]\[x(x-4)(x+3)\]\[0=x(x-4)(x+3)\]\[x=0,-3,4\](0,0) (-3,0) (4,0)

OpenStudy (anonymous):

ah okay ty! xD I forgota hah im so dumb ><

OpenStudy (anonymous):

@ChmE can you help me with another problem plsss? ><

OpenStudy (anonymous):

post the question in this feed

OpenStudy (anonymous):

well i posted it in another box so i can reward give you another medal for your helpfulness <3333